Question 9 1.81 pts How will you enter 10% in a computational formula lot in financial calculator Question 10 1.81 pts what is the future value of $12 deposited today in 5 years when earning 6% simple rate of interest? Question 11 1.81 pts In how, many years will $15O become S300 if k-02% simple interest per year? Question 12 1.81 pts At what rate of interest will an amount triple in 50 years? Enter your answer as decimals, not percentages (eg enter 001 instead of 1%). Question 13 1.81 pts Supposc an account earns a 13% simple rate of interest anually. a, The future value ot an annual deposit of S21 at the end of cach year for four yon wil be will be b. The future valuc of an annual deposit of 521 at the beginning of cach year for four years to und your final answer to two decimal places 1.81 pts Question 14
Answer #1

1) Solution: 15.60

A = P(1+rt) = 12 ( 1 + 0.06*5) = 12*1.3 = 15.60

2) Solution: 500 years

Working:

A = P(1+rt)

300 = 150 ( 1 + 0.002t)

2 = 1 + 0.002t

1 / 0.002 = t

t = 500

3) Solution: 4%

Working:

3P = P (1+50r)

3 = 1 + 50r

2 = 50r

r = 0.04 = 4%

4) Solution:

a) $100.38

b) $111.30

Working:

Year

Annual deposit

Rate

Simple interest

End value at 4 th year

1

21

0.13

8.19

29.19

2

21

0.13

5.46

26.46

3

21

0.13

2.73

23.73

4

21

0.13

0

21

100.38

Year

Annual deposit

Rate

Simple interest

End value at 4 th year

1

21

0.13

10.92

31.92

2

21

0.13

8.19

29.19

3

21

0.13

5.46

26.46

4

21

0.13

2.73

23.73

111.30